# Is Meta-training Really Necessary for Molecular Few-Shot Learning ?

### v1

@@ A.0.1 Proofs
$\max\left\{-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}\leq\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right)\leq\max\left\{-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}+\log 2.$
@@
$|\max\left\{-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}|=\min\left\{\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}\leq\frac{1}{2}\sum_{k\in\{0,1\}}\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}.$
Furthermore, $\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}=\lVert D_{k}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)\rVert^{2}_{2}$.
And $\lVert D_{k}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)\rVert^{2}_{2}\leq\lVert D_{k}\rVert_{F}^{2}\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{2}$, where $\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{2}$ is constant w.r.t ${\mathbf{M}}_{k}$,
$f_{2}\left({\mathbf{z}}_{i},\Theta\right)=O\left(\sum_{k\in\{0,1\}}\lVert D_{k}\rVert_{F}^{2}\right)=O\left(\sum_{k\in\{0,1\}}\lVert{\mathbf{M}}_{k}\rVert_{F}\right).$
Let’s study the asymptotic behavior of $\tilde{f}_{2}\left({\mathbf{z}}_{i},\Theta\right)$.
$|\tilde{f}_{2}\left({\mathbf{z}}_{i},\Theta\right)|$ $=|\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}\rvert}{\lvert\mathcal{S}\rvert}\log\det{{\mathbf{M}}_{k}}|$
$=|\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}\rvert}{\lvert\mathcal{S}\rvert}\sum_{j\in\{1\ldots d\}}\log{\sigma_{k,j}}|$
$\leq d\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}\rvert}{\lvert\mathcal{S}\rvert}\max\left\{|\log{\sigma_{k,j}}|\right\}_{j\in\{1\ldots d\}}.$
@@
As a result $\tilde{f}_{2}\left({\mathbf{z}}_{i},\Theta\right)=O\left(\sum_{k\in\{0,1\}}\log{\lVert{\mathbf{M}}_{k}\rVert_{F}}\right)$.
@@
First the log-likelihood of a gaussian distribution of mean ${\mathbf{w}}_{k}$ and of precision matrix ${\mathbf{M}}_{k}$, evaluated on ${\mathbf{z}}$ is:
$\mathcal{LL}_{{\mathbf{w}}_{k},{\mathbf{M}}_{k}}\left({\mathbf{z}}\right)=-\frac{1}{2}\left(\log\det{{\mathbf{M}}_{k}}-\lVert{\mathbf{z}}-{\mathbf{w}}_{k}\rVert_{{\mathbf{M}}_{k}}^{2}\right)+c,$
@@
$\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\left(f_{1}({\mathbf{z}}_{i},y_{i},\Theta)+\tilde{f}_{2}(\Theta)\right)$ $=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\left(\lVert{\mathbf{z}}-{\mathbf{w}}_{y_{i}}\rVert_{{\mathbf{M}}_{y_{i}}}^{2}-\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}\rvert\log\det{({\mathbf{M}}_{k})}}{\lvert\mathcal{S}\rvert}\right)$
$=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\lVert{\mathbf{z}}-{\mathbf{w}}_{y_{i}}\rVert_{{\mathbf{M}}_{y_{i}}}^{2}-\frac{1}{|\mathcal{S}|}\sum_{k\in\{0,1\}}\lvert\mathcal{S}_{k}\rvert\log\det{({\mathbf{M}}_{k})}$
$=\frac{1}{|\mathcal{S}|}\sum_{k\in\{0,1\}}\sum_{i\in\mathcal{S}_{k}}\left(\lVert{\mathbf{z}}-{\mathbf{w}}_{k}\rVert_{{\mathbf{M}}_{k}}^{2}-\log\det{({\mathbf{M}}_{k})}\right).$
For a class $k\in\{0,1\}$, we are trying to minimize the negative log-likelihood of a Gaussian distribution evaluated on the support set $\mathcal{S}_{k}$, with mean ${\mathbf{w}}_{k}$ and precision matrix ${\mathbf{M}}_{k}$.
A.1 Effect of $\lambda$ on the quadratic probe’s performances
@@
A.2 Out-of-domain experiments details
@@
A.2.1 Drug-target-interaction details
@@
A.2.2 Library screening details
@@
Results on all datasets are presented in the figure 11 and 12, as a line plot displaying the evolution of the hitrate when the number of molecules selected increases. The datasets are ordered according to the value 75% quantile of the maximum Tanimoto similarity between an active compound and the other hits in the task. This can be interpreted as a mild approximation of the task’s difficulty (as it does not consider similarity to inactive compounds or the proportion of hits).

### v2

@@ A.2 Proofs
$\max\left\{-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}\leq\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right)\leq\max\left\{-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}+\log 2.$
@@
$|\max\left\{-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}|=\min\left\{\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right\}_{k\in\{0,1\}}\leq\frac{1}{2}\sum_{k\in\{0,1\}}\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}.$
Furthermore, $\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}=\lVert D_{k}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)\rVert^{2}_{2}$.
And $\lVert D_{k}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)\rVert^{2}_{2}\leq\lVert D_{k}\rVert_{F}^{2}\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{2}$, where $\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{2}$ is constant w.r.t ${\mathbf{M}}_{k}$,
$f_{2}\left({\bm{z}}_{i},\Theta\right)=O\left(\sum_{k\in\{0,1\}}\lVert D_{k}\rVert_{F}^{2}\right)=O\left(\sum_{k\in\{0,1\}}\lVert{\mathbf{M}}_{k}\rVert_{F}\right).$
Let’s study the asymptotic behavior of $\tilde{f}_{2}\left({\bm{z}}_{i},\Theta\right)$.
$|\tilde{f}_{2}\left({\bm{z}}_{i},\Theta\right)|$ $=|\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}|}{\lvert\mathcal{S}|}\log\det{{\mathbf{M}}_{k}}|$
$=|\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}|}{\lvert\mathcal{S}|}\sum_{j\in\{1\ldots d\}}\log{\sigma_{k,j}}|$
$\leq d\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}|}{\lvert\mathcal{S}|}\max\left\{|\log{\sigma_{k,j}}|\right\}_{j\in\{1\ldots d\}}.$
@@
As a result $\tilde{f}_{2}\left({\bm{z}}_{i},\Theta\right)=O\left(\sum_{k\in\{0,1\}}\log{\lVert{\mathbf{M}}_{k}\rVert_{F}}\right)$.
@@
First the log-likelihood of a gaussian distribution of mean ${\bm{w}}_{k}$ and of precision matrix ${\mathbf{M}}_{k}$, evaluated on ${\bm{z}}$ is:
$\mathcal{L}_{{\bm{w}}_{k},{\mathbf{M}}_{k}}\left({\bm{z}}\right)=-\frac{1}{2}\left(\log\det{{\mathbf{M}}_{k}}-\lVert{\bm{z}}-{\bm{w}}_{k}\rVert_{{\mathbf{M}}_{k}}^{2}\right)+c,$
@@
$\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\left(f_{1}({\bm{z}}_{i},y_{i},\Theta)+\tilde{f}_{2}(\Theta)\right)$ $=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\left(\lVert{\bm{z}}-{\bm{w}}_{y_{i}}\rVert_{{\mathbf{M}}_{y_{i}}}^{2}-\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}|\log\det{({\mathbf{M}}_{k})}}{\lvert\mathcal{S}|}\right)$
$=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\lVert{\bm{z}}-{\bm{w}}_{y_{i}}\rVert_{{\mathbf{M}}_{y_{i}}}^{2}-\frac{1}{|\mathcal{S}|}\sum_{k\in\{0,1\}}\lvert\mathcal{S}_{k}|\log\det{({\mathbf{M}}_{k})}$
$=\frac{1}{|\mathcal{S}|}\sum_{k\in\{0,1\}}\sum_{i\in\mathcal{S}_{k}}\left(\lVert{\bm{z}}-{\bm{w}}_{k}\rVert_{{\mathbf{M}}_{k}}^{2}-\log\det{({\mathbf{M}}_{k})}\right).$
For a class $k\in\{0,1\}$, we are trying to minimize the negative log-likelihood of a Gaussian distribution evaluated on the support set $\mathcal{S}_{k}$, with mean ${\bm{w}}_{k}$ and precision matrix ${\mathbf{M}}_{k}$.
A.3 Divergence of ${\mathbf{M}}_{k}$
We study the evolution of the value of ${\mathbf{M}}_{k}$ when optimized with gradient descent (Free-Opt) and using the quadratic probe (q-probe). The results are reported in Figure 5, where we see that the maximal eigenvalue of ${\mathbf{M}}_{k}$, which diverges with gradient descent, while remaining stable using the quadratic probe (we recall that the molecule’s embedding are distributed on the unit sphere of $\mathbb{R}^{d}$).
Figure 5: Evolution of the maximal eigenvalue of ${\mathbf{M}}_{k}$ along training when optimized with gradient descent (Free-Opt) or with the quadratic probe (q-probe).
A.4 Effect of $\lambda$ on the quadratic probe’s performances
@@
A.5 Out-of-domain experiments details
@@
A.5.1 Drug-target-interaction details
@@
A.5.2 Library screening details
@@
Results on all datasets are presented in the figure 10 and 11, as a line plot displaying the evolution of the hitrate when the number of molecules selected increases. The datasets are ordered according to the value 75% quantile of the maximum Tanimoto similarity between an active compound and the other hits in the task. This can be interpreted as a mild approximation of the task’s difficulty (as it does not consider similarity to inactive compounds or the proportion of hits).
@@
A.6 Performances of other molecular pretrained backbones
Table 4: Performance comparison of pretrained with different support set sizes trained with a linear probe, compared to the linear probe applied to our multi-task backbone.
Model $|\mathcal{S}|=16$ $|\mathcal{S}|=32$ $|\mathcal{S}|=64$ $|\mathcal{S}|=128$
MolBert (Fabian et al., 2020) 0.045 $\pm$ 0.005 0.058 $\pm$ 0.005 0.074 $\pm$ 0.006 0.105 $\pm$ 0.007
AttributeMask (Sun et al., 2022) 0.048 $\pm$ 0.005 0.060 $\pm$ 0.005 0.077 $\pm$ 0.006 0.110 $\pm$ 0.007
ChemBerta-MLM (Ahmad et al., 2022) 0.049 $\pm$ 0.005 0.063 $\pm$ 0.006 0.080 $\pm$ 0.006 0.113 $\pm$ 0.007
ThreeDInfomax (Stärk et al., 2021) 0.050 $\pm$ 0.004 0.067 $\pm$ 0.005 0.081 $\pm$ 0.006 0.108 $\pm$ 0.006
ContextPred (Sun et al., 2022) 0.052 $\pm$ 0.005 0.067 $\pm$ 0.005 0.082 $\pm$ 0.006 0.113 $\pm$ 0.007
GPT-GNN (Hu et al., 2020) 0.056 $\pm$ 0.005 0.070 $\pm$ 0.005 0.085 $\pm$ 0.006 0.114 $\pm$ 0.006
MolFormer (Ross et al., 2022) 0.058 $\pm$ 0.006 0.075 $\pm$ 0.006 0.092 $\pm$ 0.007 0.125 $\pm$ 0.008
MAT (Maziarka et al., 2019) 0.052 $\pm$ 0.005 0.069 $\pm$ 0.006 0.092 $\pm$ 0.007 0.136 $\pm$ 0.009
GROVER (Rong et al., 2020) 0.068 $\pm$ 0.006 0.083 $\pm$ 0.006 0.104 $\pm$ 0.007 0.133 $\pm$ 0.007
InfoGraph (Sun et al., 2019) 0.072 $\pm$ 0.006 0.088 $\pm$ 0.006 0.106 $\pm$ 0.007 0.137 $\pm$ 0.007
GraphLog (Xu et al., 2021) 0.071 $\pm$ 0.005 0.089 $\pm$ 0.006 0.111 $\pm$ 0.006 0.138 $\pm$ 0.007
GraphCL (You et al., 2020) 0.075 $\pm$ 0.006 0.092 $\pm$ 0.006 0.112 $\pm$ 0.007 0.144 $\pm$ 0.008
GraphMVP (Liu et al., 2022) 0.078 $\pm$ 0.006 0.097 $\pm$ 0.007 0.118 $\pm$ 0.007 0.151 $\pm$ 0.008
MolR (Wang et al., 2022) 0.094 $\pm$ 0.006 0.114 $\pm$ 0.007 0.137 $\pm$ 0.007 0.170 $\pm$ 0.008
ChemBerta-MTR (Ahmad et al., 2022) 0.098 $\pm$ 0.006 0.121 $\pm$ 0.007 0.147 $\pm$ 0.008 0.183 $\pm$ 0.008
GNN-MT 0.112 $\pm$ 0.006 0.144 $\pm$ 0.007 0.177 $\pm$ 0.008 0.223 $\pm$ 0.009
CLAMP (Seidl et al., 2023) 0.202 $\pm$ 0.009 0.228 $\pm$ 0.009 0.248 $\pm$ 0.009 0.271 $\pm$ 0.009
L-probe (ours) 0.224 $\pm$ 0.010 0.252 $\pm$ 0.010 0.273 $\pm$ 0.010 0.301 $\pm$ 0.009
In the recent years, several methods were introduced proposing different ways to pre-train models in drug discovery. However, most of these models were not evaluated in a few-shot setting. We adapted several architectures on the tasks of FS-mol with the linear probe, and compare the results to the three methods already evaluated on this benchmark, that have been pretrained using the training set of FS-mol. The 14 backbones we selected for these experiments were pretrained on various datasets, such as PubChem (Wang et al., 2016), GEOM (Axelrod & Gómez-Bombarelli, 2022) and QMugs(Isert et al., 2021)), USPTO (Wang et al., 2022), and we used the weights shared publicly by the corresponding authors for each model.
We notice that none of the models enable competitive few-shot performances. This is most probably due to the fact that non of these approaches were using data originating from this benchmark in there pre-training phase. Among these models, we notice that ChemBerta-MTR (Ahmad et al., 2022) achieves the best performances, a model pretrained with a supervised multitask pretraining objective.